# Machine-learning for photoplethysmography analysis: Benchmarking feature, image, and signal-based approaches

### v1

@@ A.1.1 Blood pressure estimation
Blood pressure is one of the most widely used physiological measurements. It is a key marker of cardiovascular health; a valuable predictor of cardiovascular events; and is essential for the selection and monitoring of antihypertensive (blood pressure lowering) treatments. However, there are several limitations to blood pressure measurements currently taken in a clinical setting. First, clinical blood pressure measurements are not taken frequently in the wider population, meaning many people around the world are not aware that they have high blood pressure. Second, clinical blood pressures can be unrepresentative of a patient’s normal blood pressure due to the white coat hypertension effect, where a patient’s blood pressure is higher in the clinical setting than normal life. Third, they provide only a snapshot of blood pressure at a single time point, and do not capture diurnal variations which contain important information on health. Fourth, cuff-based blood pressure devices are not suitable for long-term monitoring. Consequently, it would be highly valuable to develop technology to monitor fluctuations in blood pressure, unobtrusively in daily life, such as in wearables.
@@
Photoplethysmography-based blood pressure estimation provides a potential approach to monitor blood pressure unobtrusively in daily life. Photoplethysmography is now widely incorporated into consumer wearable devices, and PPG signals vary with blood pressure. Researchers have focused on two main approaches to estimating blood pressure from PPG signals  [^1] Charlton PH et al., ‘Assessing hemodynamics from the photoplethysmogram to gain insights into vascular age: a review from VascAgeNet’, American Journal of Physiology-Heart and Circulatory Physiology, 322(4), H493–H522, 2022. https://doi.org/10.1152/ajpheart.00392.2021: (i) using the shape of a single PPG pulse wave; and (ii) using the speed at which the pulse wave propagates through the arteries (the pulse wave velocity). In this project, we focus on the first approach, since it can be implemented in many common wearables such as smartwatches, whereas the second requires simultaneous measurements at different anatomical sites. Blood pressure can potentially be estimated from the shape of the pulse wave using traditional signal processing based on expert-identified features, or using deep learning techniques which learn how the pulse wave is affected by blood pressure changes for themselves.
A.1.2 Atrial fibrillation detection
@@
Atrial fibrillation (AF) is the most common sustained cardiac arrhythmia and confers a five-fold increase in stroke risk. Fortunately, the risk of stroke can be greatly reduced via anticoagulation. However, patients with AF are often underdiagnosed, either because AF episodes occur asymptomatically, or because it occurs only intermittently and so is not identified during routine testing. Therefore, there is a need for continuous unobtrusive heart rhythm monitoring to identify patients with AF.
@@
Photoplethysmography provides an attractive approach to identifying AF because it is widely used in consumer wearables, and because it provides a measure of the heart rhythm, which is irregular during AF. Indeed, several smartwatches used photoplethysmography to identify irregular rhythms which may be indicative of AF [^2] Perez, M. V. et al. (2019). Large-scale assessment of a smartwatch to identify atrial fibrillation. New England Journal of Medicine, 381(20), 1909–1917. https://doi.org/10.1056/NEJMoa1901183. Irregular heart rhythms have been identified from PPG signals by identifying individual pulse waves corresponding to heartbeats, extracting inter-beat intervals, and assessing the level of inter-beat interval variability, where higher levels of variability indicate higher levels of irregularity. More recently, deep learning models have been developed to identify AF from the PPG signal  [^3] Ding, C. et al. (2024). Photoplethysmography based atrial fibrillation detection: A continually growing field. Physiological Measurement, 45(4), 04TR01. https://doi.org/10.1088/1361-6579/ad37ee.
@@ A.2 Original DeepBeat dataset
This dataset includes over 500,000 signal segments, each with a duration of 25 seconds and a sampling rate of 32 Hz, from 175 individuals, including 108 with atrial fibrillation (AF) and 67 without AF (Table LABEL:table_6). The small test set and the overlapping signal segments of the original dataset affected the analysis. This redundancy led to inflated performance metrics, while the unbalanced distribution of AF and non-AF subjects across training, validation and test sets further biased the results. Consequently, despite its impressive performance on the test set, the original model showed reduced effectiveness on validation and training data. Finally, it is important to note key limitations of this dataset for our study, all signal segments classified as AF were collected from one study protocol where patients came in for a clinical procedure. In contrast, all signal segments classified as non-AF were collected from an independent study where subjects were doing an exercise stress test. Therefore, different devices, human subject parameters (comorbidities, heart rate) would have likely influenced the classification performance, independently of AF. Therefore, confirmatory studies on independent datasets are required before any conclusions about AF classification can be drawn.
Table 5: Original DeepBeat (Stanford Wearable Photoplethysmography) Dataset
@@ A.3 Performance evaluation and metrics
Blood pressure regression In our blood pressure regression task, we use the $L_{1}$ norm to evaluate and compare the performance of the machine learning models. Specifically, we employ the empirical equivalent of this norm, known as the Mean Absolute Error (MAE), defined as follows:
$MAE\,=\,\frac{1}{N}\sum_{i=1}^{N}\left|y_{i}-\widehat{y_{i}}\right|,$ (1)
@@ A.4.1 Raw Time Series Models
Researchers have investigated a variety of deep learning models for the prediction of blood pressure using photoplethysmography (PPG) and electrocardiogram (ECG) signals. For instance, a study by Kachuee et al. (2017) [^8] Kachuee, M., Kiani, M. M., Mohammadzade, H., & Shabany, M. (2017). Cuffless blood pressure estimation algorithms for continuous health-care monitoring. IEEE Transactions on Biomedical Engineering, 64(4), 859-869. developed a deep neural network model to estimate blood pressure using ECG and PPG signals, which showed significant improvements in accuracy and reliability compared to traditional methods.
@@
Another notable study by Liang et al. (2018) [^9] Liang, Y., Chen, Z., Ward, R., & Elgendi, M. (2018). Hypertension assessment via ECG and PPG signals: An evaluation using machine learning models. Computational and Mathematical Methods in Medicine, 2018, Article 3179780 introduced a hybrid model that integrates CNN and recurrent RNN to exploit both spatial and temporal features of PPG and ECG signals. This model exhibited enhanced performance in continuous blood pressure monitoring, highlighting the potential of deep learning in medical signal processing. It is worth mentioning that transfer learning has been applied to improve the performance of blood pressure prediction models. Zhang et al. (2020) [^10] Zhang, Y., Feng, X., Li, J., Li, B., & Peng, X. (2020). Transfer learning for hybrid deep neural network-based blood pressure estimation. Journal of Healthcare Engineering, 2020, Article 5462540. utilized a pre-trained deep learning model and fine-tuned it with a smaller dataset of PPG and ECG signals. This approach minimized the requirement for extensive labelled data, enhancing the practicality of the model for real-world applications.
@@
In this project, we employed prominent models such as LeNet1d [^11] LeCun, Y., Bottou, L., Bengio, Y., & Haffner, P. (1998). Gradient-based learning applied to document recognition. Proceedings of the IEEE, 86(11), 2278-2324., Inception1d [^12] Szegedy, C., Liu, W., Jia, Y., Sermanet, P., Reed, S., Anguelov, D., … & Rabinovich, A. (2015). Going deeper with convolutions. In Proceedings of the IEEE conference on computer vision and pattern recognition (pp. 1-9)., XResNet50d, XResNet101d [^13] Strodthoff, Nils, Temesgen Mehari, Claudia Nagel, Philip J. Aston, Ashish Sundar, Claus Graff, Jørgen K. Kanters et al. ”PTB-XL+, a comprehensive electrocardiographic feature dataset.” Scientific data 10, no. 1 (2023): 279., AlexNet1d [^14] Krizhevsky, Alex, Ilya Sutskever, and Geoffrey E. Hinton. ”Imagenet classification with deep convolutional neural networks.” Advances in neural information processing systems 25 (2012). and Minirocket [^15] Dempster, Angus, Daniel F. Schmidt, and Geoffrey I. Webb. ”Minirocket: A very fast (almost) deterministic transform for time series classification.” Proceedings of the 27th ACM SIGKDD conference on knowledge discovery & data mining. 2021. as well as Temporal Convolutional Networks, to conduct a thorough assessment of the outcomes.
@@
The baseline model predicts blood pressure by using the median value of the blood pressure data, providing a straightforward reference for evaluating the performance of more advanced predictive models.
@@
The kernels do not require training, allowing for a one-time feature extraction This enables efficient model training, as only the linear regressor needs to be trained over multiple epochs. This makes the MiniRocket a very fast and effective baseline for time-series classification and regression tasks.
@@
where $x(t)$ and $y(t)$ are respectively the input and output at time t, $w_{i}$ are the weights of the convolution filter, and $k$ stands for the filter size. This equation ensures that the output $y(t)$ depends only on the current and previous inputs, preserving the causality.
@@
where $d$ represents the dilation factor (which determines the spacing between the elements of the input signal). Dilated convolutions allow the network to cover a larger range of inputs without increasing the number of parameters. Indeed, a dilated convolution lets the network look back up to $dk$ time steps, enabling exponentially larger receptive fields per the number of layers. In the original paper, the authors increased $d$ exponentially with the depth of the network.
@@ A.4.2 Clinically Interpretable Feature (CIF)-Based Models
In this study, for blood pressure estimation, PPG segments from the VitalDB dataset were preprocessed using a Butterworth infinite impulse response zero-phase band-pass filter (4th order, 0.4-7 Hz). For atrial fibrillation detection, PPG segments from the DeepBeat dataset were preprocessed using low-pass, high-pass, and adaptive filters. High-frequency noise and artifacts were removed using a low-pass infinite impulse response filter (cutoff 6 Hz), while baseline wander was eliminated using a high-pass infinite impulse response filter (cutoff 0.5 Hz) and a fifth-order normalized least mean squares adaptive filter with a reference input of unity.
@@
To estimate blood pressure (BP), 28 PPG features [^19] Charlton, P. H., Celka, P., Farukh, B., Chowienczyk, P., & Alastruey, J. (2018). Assessing mental stress from the photoplethysmogram: a numerical study. Physiological measurement, 39(5), 054001. were assessed based on pulse morphology analysis and pulse derivative features (Figure 2), which are features selected from the literature24 as the most significant for blood pressure estimation.
PPG pulse wave features include amplitude-related parameters such as the first (P1) and second (P2) systolic peaks, the diastolic peak (P3), and derived indices like the P2/P1 ratio, reflection index (RI = P3/P1), and augmentation index (AI = (P1 – P3)/P1). Time-related features consist of the pulse duration (Tp), diastolic duration (Td), systolic duration (T1), and the time interval from P1 to P3 ($\Delta$t). Area-related features include the systolic area (A1), diastolic area (A2), the inflection point area ratio (IPA = A2/A1), and the inflection point area plus the d-wave amplitude of the second PPG derivative (IPAD).
@@
To solve the blood pressure regression problem, we employed a Wavelet Packet Decomposition (WPD) approach using the Discrete Wavelet Transform (DWT) for feature extraction from raw PPG-signals. Specifically, we utilized the Daubechies wavelet (db6) of order 3, which offers a balance between time and frequency localization, making it particularly suitable for analysing non-stationary signals.
The wavelet packet decomposition allows for a hierarchical representation of the signal by decomposing both the approximation and detail coefficients at each level. This provides a finer level of signal analysis compared to standard DWT, capturing more detailed information across different frequency bands.
After obtaining the wavelet packet coefficients, these features are flattened and then directly fed, without additional processing, as inputs to a Multi-Layer Perceptron (MLP) model. This simple approach will also be used to diagnose Atrial Fibrillation from PPG-signals.
@@
Figure 3: Structure of an MLP-based atrial fibrillation detection using features extracted from the PPG signal.
@@
Figure 4: Good-quality AF PPG segment (a) with extracted PP intervals (b) and quality index (c).
Figure 5: Good-quality AF PPG segment (a) with extracted PP intervals (b) and quality index (c).
Figure 6: Bad-quality AF PPG segment (a) with extracted PP intervals (b) and quality index (c).
Figure 7: Bad-quality AF PPG segment (a) with extracted PP intervals (b) and quality index (c).
@@
  • Turning point ratio (TPR) [^21] Dash, S., Chon, K. H., Lu, S., & Raeder, E. A. (2009). Automatic real time detection of atrial fibrillation. Annals of biomedical engineering, 37, 1701-1709. evaluates randomness in PP intervals by identifying turning points—intervals greater or less than their two neighbours. It is computed as the ratio of turning points to total PP intervals within the analysis window. A higher TPR indicates greater randomness, aiding AF detection.
  • Coefficient of variation (CV) [^22] Tateno, K., & Glass, L. (2001). Automatic detection of atrial fibrillation using the coefficient of variation and density histograms of RR and $\Delta$RR intervals. Medical and Biological Engineering and Computing, 39, 664-671., [^23] Langley, P., Dewhurst, M., Di Marco, L. Y., Adams, P., Dewhurst, F., Mwita, J. C., … & Murray, A. (2012). Accuracy of algorithms for detection of atrial fibrillation from short duration beat interval recordings. Medical engineering & physics, 34(10), 1441-1447. relates the standard deviation of PP intervals to their mean. In AF, increased dispersion and decreased mean duration raise CV.
@@
Gaussian Process Regression (GPR) is a powerful and flexible non-parametric regression technique used in machine learning and statistics for modeling complex, nonlinear relationships [^32] Rasmussen, C. E. and C. K. I. Williams. Gaussian Processes for Machine Learning. MIT Press. Cambridge, Massachusetts, 2006.. Instead of fitting a specific function to the data, GPR models the relationship between input features and output as a distribution over functions. It incorporates prior knowledge (expressed through kernels) and provides uncertainty estimates for individual predictions. The objective of GPR is to find a mean function that accurately represents observed data points and can be used for predicting new data points. Gaussian Processes define a distribution over an infinite number of potential functions that could fit the data.

### v2

@@ A.1.1 BP estimation
BP is one of the most widely used physiological measurements. It is a key marker of cardiovascular health; a valuable predictor of cardiovascular events; and is essential for the selection and monitoring of antihypertensive (BP lowering) treatments. However, there are several limitations to BP measurements currently taken in a clinical setting. First, clinical BP measurements are not taken frequently in the wider population, meaning many people around the world are not aware that they have high blood pressure. Second, clinical BPs can be unrepresentative of a patient’s normal BP due to the white coat hypertension effect, where a patient’s BP is higher in the clinical setting than normal life. Third, they provide only a snapshot of blood pressure at a single time point, and do not capture diurnal variations which contain important information on health. Fourth, cuff-based blood pressure devices are not suitable for long-term monitoring. Consequently, it would be highly valuable to develop technology to monitor fluctuations in BP, unobtrusively in daily life, such as in wearables.
@@
PPG-based BP estimation provides a potential approach to monitor BP unobtrusively in daily life. PPG is now widely incorporated into consumer wearable devices, and PPG signals vary with BP. Researchers have focused on two main approaches to estimating BP from PPG signals  [^1] Charlton PH et al., ‘Assessing hemodynamics from the photoplethysmogram to gain insights into vascular age: a review from VascAgeNet’, American Journal of Physiology-Heart and Circulatory Physiology, 322(4), H493–H522, 2022. https://doi.org/10.1152/ajpheart.00392.2021: (i) using the shape of a single PPG pulse wave; and (ii) using the speed at which the pulse wave propagates through the arteries (the pulse wave velocity). In this project, we focus on the first approach, since it can be implemented in many common wearables such as smartwatches, whereas the second requires simultaneous measurements at different anatomical sites. BP can potentially be estimated from the shape of the pulse wave using traditional signal processing based on expert-identified features, or using deep learning techniques which learn how the pulse wave is affected by BP changes for themselves.
A.1.2 AF detection
@@
AF is the most common sustained cardiac arrhythmia and confers a five-fold increase in stroke risk. Fortunately, the risk of stroke can be greatly reduced via anticoagulation. However, patients with AF are often underdiagnosed, either because AF episodes occur asymptomatically or because it occurs only intermittently and so is not identified during routine testing. Therefore, there is a need for continuous unobtrusive heart rhythm monitoring to identify patients with AF.
@@
PPG provides an attractive approach to identifying AF because it is widely used in consumer wearables, and because it provides a measure of the heart rhythm, which is irregular during AF. Indeed, several smartwatches used PPG to identify irregular rhythms which may be indicative of AF [^2] Perez, M. V. et al. (2019). Large-scale assessment of a smartwatch to identify atrial fibrillation. New England Journal of Medicine, 381(20), 1909–1917. https://doi.org/10.1056/NEJMoa1901183. Irregular heart rhythms have been identified from PPG signals by identifying individual pulse waves corresponding to heartbeats, extracting inter-beat intervals, and assessing the level of inter-beat interval variability, where higher levels of variability indicate higher levels of irregularity. More recently, deep learning models have been developed to identify AF from the PPG signal  [^3] Ding, C. et al. (2024). PPG-based AF detection: A continually growing field. Physiological Measurement, 45(4), 04TR01. https://doi.org/10.1088/1361-6579/ad37ee.
@@ A.2 Original DeepBeat dataset
This dataset includes over 500,000 signal segments, each with a duration of 25 seconds and a sampling rate of 32 Hz, from 175 individuals, including 108 with AF and 67 without AF (Table LABEL:table_6). The small test set and the overlapping signal segments of the original dataset affected the analysis. This redundancy led to inflated performance metrics, while the unbalanced distribution of AF and non-AF subjects across training, validation and test sets further biased the results. Consequently, despite its impressive performance on the test set, the original model showed reduced effectiveness on validation and training data. Finally, it is important to note key limitations of this dataset for our study, all signal segments classified as AF were collected from one study protocol where patients came in for a clinical procedure. In contrast, all signal segments classified as non-AF were collected from an independent study where subjects were doing an exercise stress test. Therefore, different devices, human subject parameters (comorbidities, heart rate) would have likely influenced the classification performance, independently of AF. Therefore, confirmatory studies on independent datasets are required before any conclusions about AF classification can be drawn.
Table 9: Original DeepBeat (Stanford Wearable PPG) Dataset
@@ A.3 Performance evaluation and metrics
BP regression In our BP regression task, we use the $L_{1}$ norm to evaluate and compare the performance of the machine learning models. Specifically, we employ the empirical equivalent of this norm, known as the Mean Absolute Error (MAE), defined as follows:
[…]
@@ A.4.1 Raw Time Series Models
Researchers have investigated a variety of deep learning models for the prediction of BP using PPG and electrocardiogram (ECG) signals. For instance, a study by Kachuee et al. (2017) [^8] Kachuee, M., Kiani, M. M., Mohammadzade, H., & Shabany, M. (2017). Cuffless BP estimation algorithms for continuous health-care monitoring. IEEE Transactions on Biomedical Engineering, 64(4), 859-869. developed a deep neural network model to estimate BP using ECG and PPG signals, which showed significant improvements in accuracy and reliability compared to traditional methods.
@@
Another notable study by Liang et al. (2018) [^9] Liang, Y., Chen, Z., Ward, R., & Elgendi, M. (2018). Hypertension assessment via ECG and PPG signals: An evaluation using machine learning models. Computational and Mathematical Methods in Medicine, 2018, Article 3179780 introduced a hybrid model that integrates CNN and recurrent RNN to exploit both spatial and temporal features of PPG and ECG signals. This model exhibited enhanced performance in continuous blood pressure monitoring, highlighting the potential of deep learning in medical signal processing. It is worth mentioning that transfer learning has been applied to improve the performance of BP prediction models. Zhang et al. (2020) [^10] Zhang, Y., Feng, X., Li, J., Li, B., & Peng, X. (2020). Transfer learning for hybrid deep neural network-based BP estimation. Journal of Healthcare Engineering, 2020, Article 5462540. utilized a pre-trained deep learning model and fine-tuned it with a smaller dataset of PPG and ECG signals. This approach minimized the requirement for extensive labelled data, enhancing the practicality of the model for real-world applications.
@@
In this project, we employ prominent models such as LeNet1d [^11] LeCun, Y., Bottou, L., Bengio, Y., & Haffner, P. (1998). Gradient-based learning applied to document recognition. Proceedings of the IEEE, 86(11), 2278-2324., Inception1d [^12] Szegedy, C., Liu, W., Jia, Y., Sermanet, P., Reed, S., Anguelov, D., … & Rabinovich, A. (2015). Going deeper with convolutions. In Proceedings of the IEEE conference on computer vision and pattern recognition (pp. 1-9)., XResNet50d, XResNet101d [^13] Strodthoff, Nils, Temesgen Mehari, Claudia Nagel, Philip J. Aston, Ashish Sundar, Claus Graff, Jørgen K. Kanters et al. ”PTB-XL+, a comprehensive electrocardiographic feature dataset.” Scientific data 10, no. 1 (2023): 279., AlexNet1d [^14] Krizhevsky, Alex, Ilya Sutskever, and Geoffrey E. Hinton. ”Imagenet classification with deep convolutional neural networks.” Advances in neural information processing systems 25 (2012). and Minirocket [^15] Dempster, Angus, Daniel F. Schmidt, and Geoffrey I. Webb. ”Minirocket: A very fast (almost) deterministic transform for time series classification.” Proceedings of the 27th ACM SIGKDD conference on knowledge discovery & data mining. 2021. as well as Temporal Convolutional Networks, to conduct a thorough assessment of the outcomes.
@@
The baseline model predicts BP by using the median value of the BP data, providing a straightforward reference for evaluating the performance of more advanced predictive models.
@@
The kernels do not require training, allowing for a one-time feature extraction. This enables efficient model training, as only the linear regressor needs to be trained over multiple epochs. This makes the MiniRocket a very fast and effective baseline for time-series classification and regression tasks.
PPNet
The PPNet incorporates CNNs and LSTMs units to deal with physiological time series. CNNs characterize local spatial patterns in biosignals, and LSTMs are responsible for handling temporal dependencies, ultimately creating an efficient model to differentiate complex biomedical patterns.
iTransformer
iTransformer turns the classical transformer on its head by applying self-attention across features instead of time steps. Each variable is treated as a token, and the model can look at inter-variable interactions. Instance-normalization specificity also preserves signal characteristics, and there are improved results on multivariate time-series forecasting tasks.
TimesNet
TimesNet adopts an inception-style convolution architecture specific to time series. Through parallel convolutions with different receptive fields, it characterizes multi-periodic patterns in different scales in time.
@@
where $x(t)$ and $y(t)$ are respectively the input and output at time t, $w_{i}$ are the weights of the convolution filter, and kk\ stands for the filter size. This equation ensures that the output $y(t)$ depends only on the current and previous inputs, preserving the causality.
@@
where dd\ represents the dilation factor (which determines the spacing between the elements of the input signal). Dilated convolutions allow the network to cover a larger range of inputs without increasing the number of parameters. Indeed, a dilated convolution lets the network look back up to $dk$ time steps, enabling exponentially larger receptive fields per the number of layers. In the original paper, the authors increased $d$ exponentially with the depth of the network.
@@ A.4.2 Clinically Interpretable Feature (CIF)-Based Models
In this study, for BP estimation, PPG segments from the VitalDB dataset were preprocessed using a Butterworth infinite impulse response zero-phase band-pass filter (4th order, 0.4-7 Hz). For atrial fibrillation detection, PPG segments from the DeepBeat dataset were preprocessed using low-pass, high-pass, and adaptive filters. High-frequency noise and artifacts were removed using a low-pass infinite impulse response filter (cutoff 6 Hz), while baseline wander was eliminated using a high-pass infinite impulse response filter (cutoff 0.5 Hz) and a fifth-order normalized least mean squares adaptive filter with a reference input of unity.
@@
To estimate BP, 28 PPG features [^19] Charlton, P. H., Celka, P., Farukh, B., Chowienczyk, P., & Alastruey, J. (2018). Assessing mental stress from the photoplethysmogram: a numerical study. Physiological measurement, 39(5), 054001. were assessed based on pulse morphology analysis and pulse derivative features (Figure 2), which are features selected from the literature24 as the most significant for BP estimation.
[…]
@@
To solve the BP regression problem, we employ a Wavelet Packet Decomposition (WPD) approach using the Discrete Wavelet Transform (DWT) for feature extraction from raw PPG-signals. Specifically, we utilize the Daubechies wavelet (db6) of order 3, which offers a balance between time and frequency localization, making it particularly suitable for analysing non-stationary signals.
The wavelet packet decomposition allows for a hierarchical representation of the signal by decomposing both the approximation and detail the coefficients at each level. This provides a finer level of signal analysis compared to standard DWT, capturing more detailed information across different frequency bands.
After obtaining the wavelet packet coefficients, these features are flattened and then directly fed, without additional processing, as inputs to an MLP model. This simple approach will also be used to diagnose AF from PPG-signals.
@@
Figure 3: Structure of an MLP-based model for the AF detection task using features extracted from the PPG signal.
@@
Figure 4: Good-quality PPG segment for the AF task (a) with extracted PP intervals (b) and quality index (c).
Figure 5: Good-quality PPG segment for the AF task (a) with extracted PP intervals (b) and quality index (c).
Figure 6: Bad-quality PPG segment for the AF task (a) with extracted PP intervals (b) and quality index (c).
Figure 7: Bad-quality PPG segment for the AF task (a) with extracted PP intervals (b) and quality index (c).
@@
  • Turning point ratio (TPR) [^22] Dash, S., Chon, K. H., Lu, S., & Raeder, E. A. (2009). Automatic real-time detection of atrial fibrillation. Annals of biomedical engineering, 37, 1701-1709. evaluates randomness in PP intervals by identifying turning points—intervals greater or less than their two neighbours. It is computed as the ratio of turning points to total PP intervals within the analysis window. A higher TPR indicates greater randomness, aiding AF detection.
[…]
@@
Gaussian Process Regression (GPR) is a powerful and flexible non-parametric regression technique used in machine learning and statistics for modelling complex, nonlinear relationships [^33] Rasmussen, C. E. and C. K. I. Williams. Gaussian Processes for Machine Learning. MIT Press. Cambridge, Massachusetts, 2006.. Instead of fitting a specific function to the data, GPR models the relationship between input features and output as a distribution over functions. It incorporates prior knowledge (expressed through kernels) and provides uncertainty estimates for individual predictions. The objective of GPR is to find a mean function that accurately represents observed data points and can be used for predicting new data points. Gaussian Processes define a distribution over an infinite number of potential functions that could fit the data.